OpenStudy (anonymous):

If f(x) = 0.8(2 – x), what is the value of f(–2)? A. 0 B. 1.6 C. 3.2 D. 3.6

OpenStudy (anonymous):

3hunna Bang!!!!!!!

OpenStudy (anonymous):

what

OpenStudy (anonymous):

f(-2) means x= -2 So everywhere you see an X, you plus in -2 So it would be: 0.8[2-(-2)] Simplify: 0.8(0) Answer: 0

OpenStudy (anonymous):

* plug in

OpenStudy (anonymous):

C

OpenStudy (anonymous):

f(x) = 0.8(2 - x) f(x) = 0.8(2 - (-2)) f(x) = 0.8(2 + 2) f(x) = 0.8(4) f(x) = 3.2

OpenStudy (anonymous):

@WatchDogs

OpenStudy (anonymous):

@PikaPaige6 How does [2 - (-2)] = 0?

OpenStudy (anonymous):

A negative and a negative = a positive. So it becomes 2 + 2.

OpenStudy (anonymous):

Oh oh I see that now. I must have worked it out too fast. Thank you.

OpenStudy (anonymous):

If it was [2 - (+2)] it would = 0. But in this case they're both negatives, and it equals 4.

OpenStudy (anonymous):

No problem.